### v3

# GigaHands: A Massive Annotated Dataset of Bimanual Hand Activities

Abstract
Understanding bimanual human hand activities is a critical problem in AI and robotics. We cannot build large models of bimanual activities because existing datasets lack the scale, coverage of diverse hand activities, and detailed annotations. We introduce GigaHands, a massive annotated dataset capturing 34 hours of bimanual hand activities from 56 subjects and 417 objects, totaling 14k motion clips derived from 183 million frames paired with 84k text annotations. Our markerless capture setup and data acquisition protocol enable fully automatic 3D hand and object estimation while minimizing the effort required for text annotation. The scale and diversity of GigaHands enable broad applications, including text-driven action synthesis, hand motion captioning, and dynamic radiance field reconstruction.
Figure 1: GigaHands is a massive dataset of human bimanual activities with paired text annotations. Each column above shows an activity sequence from the dataset. The dataset covers diverse 3D hand activities, including hand-object interactions (blue) across object scales, gestures (orange), and self-interactions (red). Each clip is paired with descriptive text and 51 camera views, enabling radiance field reconstruction. The bottom row show other annotations in the dataset including hand shape, object shape and pose (left half images). The right half images show novel views from dynamic radiance field fitting.
1 Introduction
“The human hand is a marvel of evolution, whose intricate structures and capabilities have allowed us to manipulate the environment in ways no other species can.”
– F.R. Wilson [wilson1999hand]
From the skillful manipulation involved in cooking a meal, to the rapid movement of fingers to type this sentence, hands are always busy shaping our environments and communicating with others. Enabling machines to understand and replicate the remarkable capabilities of human hands is one of the grand challenges of AI and robotics. One approach to achieve this goal could be to train models on massive datasets, the strategy followed by recent large models that learn from trillions of text tokens [dubey2024llama], billions of text-image pairs [schuhmann2022laion], tens of millions of 3D shapes [deitke2024objaverse], or millions of robot trajectories [o2023open, brohan2022rt, brohan2023rt, kim2024openvla]. However, datasets of this scale are currently unavailable for natural hand activities.
Sourcing large-scale 3D hand activities data is challenging. The two most common methods for data acquisition involve using cameras to capture hand manipulations in the wild or in controlled studio settings. In-the-wild data includes monocular internet videos [shan2020understanding], egocentric videos captured using wearable cameras [pirsiavash2012detecting, damen2018scaling, li2021eye, damen2022rescaling, grauman2022ego4d, liu2022hoi4d], or multi-view videos from third-person cameras [goyal2017something, miech2019howto100m, shan2020understanding]. However, this data is sparse, hard to calibrate, and noisy, resulting in limited 3D motion reconstruction accuracy, especially for objects. Alternatively, studio settings bring subjects into camera-rich environments and could employ markers for accurate reconstruction. However, the staged setup and lack of real-world context limits data diversity, and marker-based tracking [liu2022beat, garcia2018first, taheri2020grab, yang2022oakink, fan2023arctic, liu2024taco, zhan2024oakink2, banerjee2024hot3d, lv2025himo] inhibits natural interactions.
[…]
GigaHands can unlock new capabilities in applications including motion generation, robotics, and dynamic 3D reconstruction [pumarola2021d]. To demonstrate, we show improvements enabled by our dataset on text-driven action synthesis and hand motion captioning, along with examples dynamic radiance field reconstruction. To sum up our contributions:
  • We present GigaHands, a massive, diverse and annotated 3D bimanual activities dataset. It includes 34 hours of activities, 14k hand motions clips paired with 84k text annotation, and over 183M unique hand images.
  • A procedural data acquisition strategy that ensures activity diversity and detailed atomic-level text descriptions while minimizing manual effort. We also built an accurate and fully-automated pipeline for estimating 3D hand/object shape and pose, segmentation, and camera pose.
  • We show applications enabled by our dataset’s scale, including 3D hand motion generation, hand motion captioning, and dynamic semantic scene reconstruction.
2 Related Work
2.1 Hand Motion Data Sourcing
Hand motion data has four primary sources, and each presents its own advantages and challenges concerning realism, diversity, accuracy, and practicality.
Static poses involve motion planning [amor2012generalization, bai2014dexterous, brahmbhatt2019contactgrasp, she2022learning], policy learning [mandikal2021learning, christen2022d], or motion synthesis [zhang2024artigrasp, christen2024diffh2o, zhang2024graspxl] from individual static positions. While these methods provide reliable 3D data, it lacks semantic richness, resulting in motions disconnected from real-world activities and not fully representing diverse human movement.
Synthetic data from game engines [zimmermann2017learning], augmented reality [mueller2017real], virtual reality [han2022umetrack], or simulations [miller2004graspit, hasson2019learning, turpin2022grasp] allow for controlled and accurate 3D motion sequences. Despite their reliability, these synthetic motions often fail to capture the intricacies of genuine human movements, limiting their applicability in modeling natural hand behaviors.
Real-world in-the-wild settings involves using wearable or portable sensors to capture hand interactions from egocentric [pirsiavash2012detecting, damen2018scaling, li2021eye, damen2022rescaling, grauman2022ego4d, liu2022hoi4d], third-person [goyal2017something, miech2019howto100m, shan2020understanding] or both types of views [sigurdsson2018charades, grauman2024ego], yielding realistic and natural motions. However, obtaining reliable 3D motion data for both hands and objects is challenging due to occlusions resulting from limited sensor deployment and accuracy. Additionally, annotating such data is labor-intensive, and certain self-contact motions are difficult to record accurately.
Studio environments place participants in controlled settings equipped with extensive sensors, enabling reliable capture of detailed information such as 3D hand motions [joo2018total, zhang2022egobody, liu2022beat, ohkawa2023assemblyhands, lin2024motion, zhang2024both2hands], object movements [garcia2018first, hampali2020honnotate, chao2021dexycb, kwon2021h2o, hampali2022keypoint, yang2022oakink, banerjee2024hot3d, liu2024taco, zhan2024oakink2, lv2025himo], contact regions [taheri2020grab, fan2023arctic, pokhariya2024manus], audio [lee2019talking, yoon2022genea, jin2024audio, ng2024audio], and tactile data [buscher2015flexible, pham2017hand, sundaram2019learning, grady2022pressurevision]. However, collecting data in studios is arduous and may not reflect the diversity of real-life scenarios. Participants might find it challenging to perform natural motions in an unfamiliar environment, especially when encumbered with wearable sensors. For example, motion capture (MoCap) systems [liu2022beat, garcia2018first, taheri2020grab, yang2022oakink, fan2023arctic, liu2024taco, zhan2024oakink2, banerjee2024hot3d, lv2025himo] require markers attached to the body, reducing visual realism and inhibiting movements like self-contact between hands, necessitating additional post-processing to restore a realistic appearance. On the other hand, markerless motion capture systems [joo2018total, hampali2020honnotate, chao2021dexycb, kwon2021h2o, hampali2022keypoint, zhang2022egobody, ohkawa2023assemblyhands, lin2024motion, pokhariya2024manus, zhang2024both2hands] offer more realistic visuals and encourage natural motions, but they may trade off capture accuracy. GigaHands balances accuracy, diversity, realism, and practicality by replicating in-the-wild settings during activity elicitation and estimating accurate 3D motion from marker-less motion capture.
2.2 Hand Datasets and Annotations
Hand datasets vary in sources and scales, each constructed for specific purposes and enriched with various annotations.
Hand motion is the most common type of annotation, represented as bounding boxes [pirsiavash2012detecting, goyal2017something, shan2020understanding], segmentation masks [narasimhaswamy2019contextual, li2021eye, darkhalil2022epic], 2D keypoints [andriluka20142d, mckeenz, jin2020whole, sener2022assembly101], 3D keypoints [qian2014realtime, tang2014latent, sun2015cascaded, tzionas2016capturing], or parameters of parametric models [sridhar2013interactive, liu2022beat, pavlakos2024reconstructing]. Depending on the data source and desired detail, annotations are obtained through manual labeling [tzionas2016capturing], synthesis [sharp2015accurate, zimmermann2017learning, mueller2017real, mueller2018ganerated], markers and sensors [tompson2014real, yuan2017bighand2, garcia2018first, joo2018total, zhang2022egobody, lin2024motion, zhang2024both2hands], cross-view bootstrapping [simon2017hand, moon2020interhand2], iteratively trained keypoint extraction networks [zimmermann2019freihand, ohkawa2023assemblyhands], multi-view annotation [ohkawa2023assemblyhands], or audio refinement [jin2024audio]. Semi-automatic and automatic labeling techniques significantly improve scalability.
Object annotations are provided given the frequent interaction between hands and objects. These can be acquired from manual labeling [sridhar2016real, chao2021dexycb, liu2022hoi4d], synthesis [hasson2019learning, corona2020ganhand], MoCap [ye2021h2o, yang2022oakink, fan2023arctic, banerjee2024hot3d, liu2024taco, zhan2024oakink2, zhao2024m, lv2025himo], hand-object reconstruction or retrieval [cao2021reconstructing, xie2023hmdo, min2024genheld], or multi-view RGB-D data [hampali2020honnotate, kwon2021h2o, hampali2022keypoint]. We demonstrate that multi-view RGB us sufficient for automatic annotation, enhancing scalability.
Text annotations. To capture the rich semantic meanings in human hand actions, certain datasets provide text annotations. These annotations may include action types [pirsiavash2012detecting, goyal2017something, dreher2019learning, krebs2021kit, grauman2022ego4d], atomic action descriptions [damen2018scaling, li2021eye, damen2022rescaling, sener2022assembly101, grauman2024ego], activity narration [damen2018scaling, miech2019howto100m, damen2022rescaling, grauman2024ego], activity commentary [grauman2022ego4d, grauman2024ego, zhan2024oakink2, cha2024text2hoi], object affordance [yang2022oakink, jian2023affordpose, zhan2024oakink2, liu2024taco], or body dynamics [zhang2024nl2contact, zhang2024both2hands, yu2024signavatarslargescale3dsign]. Typically, annotations are manually provided, especially when data are collected in uncontrolled, unscripted environments [damen2018scaling, damen2022rescaling, grauman2022ego4d, grauman2024ego]. Even in studio settings, post-processing is often required to extract meaningful action clips [li2021eye, zhan2024oakink2]. We introduce a procedural instruct-and-annotate pipeline that maximizes semantic interpretability and minimizes the annotation effort.
Contact are derived from manual annotation [shan2020understanding], video grounding [nagarajan2019grounded], thermal sensors [brahmbhatt2019contactdb, brahmbhatt2020contactpose] or geometry analysis [taheri2020grab, zhu2024contactart, pokhariya2024manus] based on captured data. Our dataset enables us to derive contact regions using geometry analysis on both surface mesh and density fields.
Table 1: Comparisons of 3D bimanual motion datasets. Dataset names are highlighted with different colors if it has no text annotations (gray), action type (green), sparse description (red), and dense description (blue). See the supp. document for the full table.
Name mins motions poses views frames sub. obj.
w/o Object Annotation
AssemblyHands [ohkawa2023assemblyhands] 630 62 203k 12 3.03M 34 ///
Ego-Exo4D [grauman2024ego] /// /// 4.4M 5-6 /// 740 ///
w/ Object Annotation
HOI4D [liu2022hoi4d] 1,333 4k 1.2M 1 2.4M 4 800
ARCTIC [fan2023arctic] 121 339 218k 9 2.1M 10 11
TACO [liu2024taco] 202 2.3k 363k 13 4.7M 14 196
OakInk2 [zhan2024oakink2] 557 2.8k 993k 4 4.01M 9 75
HOT3D [banerjee2024hot3d] 833 4.1k 1.7M 2-3 3.7M 19 33
GigaHands (Ours) 2,034 13.9k 3.7M 51 183M 56 417
Figure 2: Dataset Diversity. The left and middle figures illustrate the diversity of pose and motion variations in GigaHands, visualized using t-SNE embeddings. Some points along the convex hull are highlighted with their corresponding text instructions, showcasing unique motions captured in our dataset. The right figure compares the verb sets among different datasets using an UpSet visualization [lex2014upset]. Each column represents the number of verbs exclusive to specific subsets of datasets, indicated by the connected dots below the columns. The rows indicate the total verb count in each dataset. GigaHands contains more verbs and more exclusive verbs compared to other datasets.
3 The GigaHands Dataset
Dataset Characteristics.
GigaHands is a large and diverse dataset of bimanual hand activities with detailed annotations. It encompasses a wide range of scenarios, including hand-object interactions, gesturing, and self-contacts performed by 56 subjects who collectively used 417 objects (see Table 1). The dataset contains 2,034 minutes of bimanual hand activities, surpassing the length of any existing 3D hand motion dataset. Compared to other datasets that are unannotated [fan2023arctic, banerjee2024hot3d], contain motion type annotations [ohkawa2023assemblyhands, liu2022hoi4d, liu2024taco], or provide sparse motion descriptions [grauman2024ego, zhan2024oakink2], GigaHands offers detailed text annotations for all captured activities. With a total of 13k instructed motion sequences, 14k annotated motion clips, and 84k augmented text descriptions, it is larger than any other dataset. Furthermore, GigaHands includes 3.7 million bimanual 3D hand poses, represented by both 3D keypoints and MANO [romero2022embodied] hand meshes, comparable to the scale of Ego-Exo4D [grauman2024ego] – the current largest annotated 3D hand pose dataset. However, unlike Ego-Exo4D, which is annotated from selected, non-continuous frames, GigaHands provides continuous bimanual 3D hand poses and shapes. Each motion clip was captured using 51 camera views, resulting in 183M RGB frames (and 366M unique hand images). This multi-view setup enables new applications such as dynamic 3D radiance field reconstruction. Additionally, the dataset contains annotations for hand/object segmentation masks, 3D object shape, pose and appearance, 2D/3D keypoints, and camera poses.
Diverse Hand Pose and Motion.
Figure 2 (left) compares hand pose diversity across datasets using t-SNE [van2008visualizing] clustering of 3D keypoints of both hands. Similarly, Figure 2 (middle) shows hand motion diversity through t-SNE clustering of the latent codes of a Variational Autoencoder trained for motion reconstruction. GigaHands exhibits significantly greater diversity than existing datasets in both bimanual hand pose and motion. Additionally, datapoints near the boundary of our dataset seem to correspond to text annotations that are unique in our datasets, demonstrating that diverse verbs contribute to diverse poses and motions.
Diverse Text Annotation.
GigaHands contains activities that are more diverse than existing datasets [zhan2024oakink2, grauman2022ego4d, grauman2024ego] as measured using verb counts in Figure 2 (right). Since each verb corresponds to an activity, a diverse set of verb indicates activity diversity. GigaHands contains the most verbs (1467) with 580 of them being unique to our dataset. The key to achieving this diversity is our instruct-to-annotate strategy (see Section 4.3) for automated labeled activity sourcing. Please refer to the supp. document for the source of verbs and detailed verb comparison.
Figure 3: Diverse Objects and Frequent Hand Contact Regions. GigaHands provide objects (left) spanning diverse scenarios, including cooking, office working, crafting, entertainment, and housework. The diverse activities result in contact regions (right) spanning both the front and back of both hands.
Figure 4: Instruct-to-Annotate Pipeline. The instruction elicitation process (left yellow block) creates atomic action-level instruction scripts in a temporally smooth order, structured within scenes. This is achieved by parsing action datasets, grouping verbs into a pool, structuring scenarios, and generating scene scripts. During filming, subjects act according to these scripts, producing recorded motion sequences. Annotators then process these sequences (right blue block) by segmenting them into clips and annotating unscripted motions.
Diverse Objects.
Figure 3 shows some of the tabletop objects included in GigaHands. We provide 3D meshes [arcodeCodeAugmented] with associated textures for rigid objects and multi-view segmentation masks for non-rigid objects. Different from most datasets, we also include small-scale objects (e.g. pens) that are difficult to capture. We employ single-view reconstruction [meshyMeshyDocs] to obtain meshes for these small objects. In total, the dataset includes 417 objects, with 310 multi-view scanned meshes and 31 single-view generated meshes.
Diverse Contacts.
Since we have 3D meshes both hands and objects, we can use them to estimate contact maps on both hands, following the approach in [taheri2020grab]. Figure 3 shows the accumulated contact regions from randomly sampled frames across the dataset. The results reveal diverse contact areas, including regions between fingers and on the back of the hands (e.g. punching uses the backs of hands). The right hand interacts with objects more frequently than the left, since a majority of our subjects were right-handed.
4 Dataset Acquisition
We present our data acquisition pipeline, which balances realism, diversity, and accuracy while reducing annotation effort. Our procedural pipeline, called Instruct-to-Annotate (see Figure 4), consists of several key components:
  • Procedural instruction elicitation (Section 4.1) for generating instructions to guide participants during data capture.
  • Filming (Section 4.2) using a markerless multi-camera system to record high-quality hand activity sequences.
  • Motion annotation and text augmentation (Section 4.3) to refine sequences by breaking them into shorter clips with accurate annotations.
  • Hand motion (Section 4.4) and object motion estimation (Section 4.5) for detailed 3D hand and object motions from the captured multi-view RGB videos.
4.1 Instruction Elicitation
To ensure diverse hand activities and reduce annotation effort, the Instruct-to-Annotate pipeline starts with a procedural instruction elicitation protocol. We began by sourcing atomic actions from multiple datasets [grauman2022ego4d, grauman2024ego, zhan2024oakink2, liu2024taco], extracting a pool of verbs corresponding to those actions. To elicit subjects to perform these verbs (actions), we manually associated each verb with multiple objects and, with the assistance of an LLM [achiam2023gpt], grouped the verbs and objects into different scenarios such as cooking and eating, office work, crafting, entertainment, and housework. We then structured these scenarios into scenes where the objects could co-occur. Using LLM, we organized the scenes into lists of activities that utilized the objects and verbs in a temporally smooth order and automatically generated detailed instruction scripts. These scripts comprise 5 scenarios, 25 scenes, 191 activities, and 1370 instructions containing a total of 533 verbs. Please see the supp. document for details.
4.2 Filming
Hardware.
[…]
Filming Process.
During filming, subjects perform actions according to the provided instructions, keeping both hands within the filming area. Instructions are converted to audio and played sequentially to guide the participants, while an operator controls the capture by playing each audio instruction and recording each motion sequence. If the ending state of a performance does not align with the next instruction, a corrective instruction will be given and re-recorded to ensure smooth transitions and reduce annotation efforts. This approach ensures all recorded sequences correspond to a pre-scripted or recorded instruction.
4.3 Action Annotation & Augmentation
Though each filmed motion sequence was paired with an instruction, manual annotation was still necessary for two reasons. First, the instructions generated by the LLM occasionally contained inconsistencies or hallucinations. Second, participants might misunderstand the instructions or add unintentional actions during filming. To encourage subjects to act freely and naturally, we intentionally retained these recordings. Annotators then split these sequences into individual clips and annotated any actions not included in the original instructions. This process resulted in a more accurate dataset with matched (description, clip) pairs. In total, we refined the 13k motion sequences into 14k motion clips.
To further enhance the text descriptions, we used the LLM to rephrase each description 5 times, providing multiple textual variations for each motion clip. This expanded our 14k motion clips into 84k motion-text pairs containing 1,467 unique verbs. The augmentation prompts and examples are provided in the supp. document.
4.4 Hand Motion Estimation
GigaHands provides detailed hand motion data, including 2D and 3D hand keypoints and MANO [romero2022embodied] meshes for both hands. Since existing hand shape and pose estimation did not work well enough, we built our own hybrid method. We begin by obtaining bounding boxes for the hands across videos using YOLOv8 [reis2023real]. Then, 2D keypoints are extracted from the MANO meshes estimated with HaMeR [pavlakos2024reconstructing] and handedness (left or right) is determined with ViTPose [xu2022vitpose]. HaMeR meshes cannot directly be used since they lacks accurate depth. With camera parameters, 2D keypoints are triangulated across views [sridhar2013interactive] to obtain accurate 3D positions. To ensure temporal smoothness in the sequences, the one-euro filter [casiez20121] is applied to the both 2D and 3D keypoints. With the bounding box and 2D & 3D keypoints, we fit the MANO parametric hand model [romero2022embodied] following the EasyMoCap pipeline [easymocap]. This results in a fully automated pipeline for extracting coherent and accurate hand motions. Details and the evaluation of each step are provided in the supp. document.
4.5 Object Motion Estimation
GigaHands also provides 3D object motions represented as 3D shape and 6D pose. Given target meshes obtained from pre-scanning or single-view reconstruction, we track these objects with multi-view constraints. First, we segment the target objects from the background in each view. We detect salient objects across the video using DINOv2 [oquab2023dinov2] on frames subsampled at 1 fps. Using both object text descriptions and rendered template meshes from multiple views, we select the top-k bounding boxes most aligned with the template mesh throughout the video with Grounding DINO [liu2023grounding]. To eliminate false positives, we use OpenCLIP [ilharco_gabriel_2021_5143773] to filter out boxes aligned with negative prompts. With the positive bounding boxes as the prompt, we use SAM2 [ravi2024sam] to segment object masks throughout the video.
Since no existing method for object pose estimation worked well [ornek2023foundpose], we decided to build our own robust method that exploits our dense multi-view setting. We use a differentiable rendering approach [ravi2020pytorch3d] supervised by the multi-view masks. To initialize the translation and rotation, we first build a radiance field using Instant-NGP [muller2022instant]. We initialize the object’s translation using the center of the density field. To address potential object symmetries, we use object appearance to initialize the rotation. Following FoundPose [ornek2023foundpose], we render the template mesh with multiple initial rotations and use DINOv2 features to find the best match, aggregating cross-view information. This process yields precise object motion estimates even in complex, cluttered scenes. More details and the evaluation of each step are provided in the supp. document.
5 Applications & Experiments
We demonstrate the utility of GigaHands in several applications that require large-scale data, including text-driven hand motion synthesis (Section 5.1), motion captioning for our dataset (Section 5.2) and for in-the-wild datasets (Section 5.2), and dynamic radiance field reconstruction (Section 5.3). More results can be found in the supp. document.
Figure 5: Generated motions from models trained on different datasets. Texts highlighted in green, orange, and blue come from the OakInk2, TACO, and GigaHands datasets. In the bottom two rows, hand meshes highlighted in these colors are generated by models trained on the corresponding datasets. The model trained on GigaHands can generate diverse motions from a single text (right four columns) and accurate motion with text from other datasets (left two columns). Darker color indicates later frame in the sequence.
5.1 Text-driven Hand Motion Synthesis
Generating diverse and complex motions is crucial for training virtual agents and robotic manipulation. We use GigaHands to train models for text-driven hand motion synthesis, overcoming the limitations of approaches that rely on strict conditions [zhan2024oakink2] or generate only simple skills [cha2024text2hoi] due to constrained data. We demonstrate that the scale of GigaHands leads to better performance compared to other datasets [liu2024taco, zhan2024oakink2] and smaller subsets of GigaHands.
Training and Evaluation Protocol.
We split GigaHands into train, test, and val sets with a ratio of 16:3:1. From among our annotations, we chose $4242$ 3D keypoints for both hands as the representation for model training (see the supp. document for an analysis of alternative representations). To evaluate the naturalness, diversity, and alignment of generated motions with textual descriptions, we use metrics from [Guo_2022_CVPR], including R-Precision, Multimodal Distance (MM Dist), Fréchet Inception Distance (FID), Diversity (Div.), and Multimodality (MM.). For computing FID and Div., we train a motion autoencoder to define a compressed motion space. For R-Precision, MM Dist, and MM., we employ contrastive learning to create a joint motion-text embedding space, following [Guo_2022_CVPR]. Feature extractors are trained independently for each dataset and subset.
Table 2: Quantitative results for text-driven motion synthesis with models trained on different datasets. upper bound indicates performance calculated with the ground truth. We report the mean of 20 evaluations, and $\rightarrow$ means the closer to the upper bound the better. The model trained on GigaHands performs best on most metrics.
[…]
@1 @2 @3
upper bound 64.4 86.2 89.4 2.86 0.045 14.2 -
TACO [liu2024taco] 18.9 37.7 52.9 7.39 11.0 11.1 6.83
upper bound 50.4 71.2 81.1 3.67 0.022 9.30 -
OakInk2 [zhan2024oakink2] 17.9 31.7 47.9 7.75 19.6 6.88 3.45
upper bound 77.4 88.8 91.3 2.96 0.002 11.9 -
GigaHands 31.2 44.7 53.1 6.68 4.70 10.5 9.11
Results.
We report text to hand motion synthesis performance using the T2M-GPT [zhang2023generating] backbone in Table 2 (see the supp. document for more results on GRU-based [chuan2022tm2t] and diffusion [tevet2023human] architectures). Since the evaluation embedding spaces are trained on different test sets, we also evaluate the metrics on ground truth test data to indicate upper-bound performance for each test set. Models trained on GigaHands outperform others on all metrics except MM Dist., with higher R-Precision due to better text-motion alignment. Compared to OakInk2 (which also includes textual descriptions for hand motions), GigaHands achieves significantly better FID, Diversity, and Multimodality scores, likely due to its greater dataset diversity. Figure 5 shows hand motions generated by T2M-GPT trained with GigaHands using text inputs from the test sets of GigaHands, OakInk and TACO. Even without object geometry input, our model generates reasonable hand shapes and poses for object manipulation, benefiting from the diversity of objects in our dataset. Our model can also generate reasonable motion using text from other datsasets, showing the comprehensiveness of GigaHands.
Figure 6: Effect of dataset size on motion reconstruction and text-to-motion generation performance. The x-axis shows the percentage of training data used (10%, 20%, 50%, 80%, and 100%), and the y-axis displays performance metrics: FID, MM Dist., Top-1, and Top-3 accuracy. Larger datasets consistently improve performance across all metrics, highlighting the benefits of increased data scale.
Effect of Data Scale.
Figure 6 illustrates the impact of dataset size on motion reconstruction and text-to-motion generation. Based on the T2M-GPT architecture, we train a motion VQ-VAE for reconstruction and a generative pretrained transformer model for text-to-motion generation using 10%, 20%, 50%, 80%, and 100% of the training set, evaluating on the same test set. We evaluate FID, MM Dist., and Top 1 & Top 3 accuracies. Most metrics continually improve with larger datasets, demonstrating the value of larger-scale data.
5.2 Hand Motion Captioning
Captioning human hand motion helps interpret action intent. We train a motion captioning model on GigaHands and generate captions for its test set motions; we also show examples of captioning in-the-wild datasets.
Training and Evaluation Protocol.
We adopt TM2T [chuan2022tm2t] as the model backbone. For text-motion alignment evaluation, we utilize R-Precision and MM Dist., employing the same embedding space described in Section 5.1. To evaluate alignment with the ground-truth text annotations, we use BLEU [papineni2002bleu], ROUGE [lin2004rouge], and BERTScore [zhang2019bertscore]. Since our dataset has diverse verbs, we evaluate text diversity with distinct-n [li2015diversity] and Pairwise BLEU [shen2019mixture]. Since TACO lacks text annotations, we use its triplet labels (<action, tool, object>) as scripts, transforming motion captioning in TACO into a classification task.
Table 3: Quantitative evaluation for motion captioning with models trained on different datasets. GigaHands performs best on most metrics.
[…]
upper bound 64.4 86.2 94.2 2.86
TACO [liu2024taco] 43.5 59.6 68.5 5.84
upper bound 51.1 69.5 79.6 3.67
Oakink2 [zhan2024oakink2] 40.4 58.8 68.2 4.55
upper bound 75.3 89.1 93.9 2.87
GigaHands 57.0 66.1 69.8 5.37
Table 4: Pairewise-BLEU, BLEU@4, ROUGE, distinct-n, and BERTScore for motion captioning with models trained on different datasets. GigaHands performs best on most metrics.
[…]
TACO [liu2024taco] 4.59 39.4 61.2 2.03 6.35 57.1
Oakink2 [zhan2024oakink2] 8.21 39.9 56.3 2.76 6.82 35.3
GigaHands 0.916 43.1 57.7 15.3 36.9 55.4
Figure 7: Motion captioning results with different datasets. Each column shows a motion sequence, its ground truth text description, and two generated texts. Hand motions highlighted in green, orange, and blue come from OakInk2, TACO, and GigaHands, respectively. Texts highlighted in these colors are generated by models trained on the corresponding datasets. The model trained on GigaHands generates diverse captions from a single motion (right three columns) and accurately captions motions from other datasets (left two columns).
Figure 8: Synthesized test views using 2DGS for the motion ‘zip up the pants,’ displayed at two timesteps and from two viewpoints. The synthesized views faithfully reconstruct the scene by leveraging the ample camera views provided by GigaHands.
Hand Motion Captioning for GigaHands
Figure 7 shows examples of generated 3D motion captions. The model trained on GigaHands generates diverse and accurate captions. Though objects are not present in the input, the model still generates captions which include reasonable object descriptions. Table 3 and Table 4 report the quantitative performance of models trained on different datasets [zhan2024oakink2, liu2024taco]. The model trained on GigaHands achieves superior motion-text alignment, particularly in Pairwise BLEU and distinct-n, which reflect caption diversity. While our accuracy is marginally lower compared to the simple triplet captioning in TACO, our approach produces richer and more varied captions.
3D Hand Motion Captioning for Unlabeled Datasets
GigaHands’s large variety motions and associated annotations allows models trained on it to caption other 3D hand motion datasets. Fig. 7 shows some examples using the model trained on GigaHands to caption motions from TACO and OakInk2 datasets after aligning their motion range with GigaHands. This model trained with GigaHands generates finer-granularity captions for these other datasets.
5.3 Dynamic Radiance Field Reconstruction
Unlike existing hand activity datasets, GigaHands provides dense camera views. This enables new applications, such as building dynamic radiance fields. From the 51 views provided, we remove 12 camera views due to lighting issues and randomly select one view as the test view, resulting in 38 views used for training. We segment consistent object and hand masks throughout the video as described in Section 4.5. We then fit 2DGS [huang20242d] for frame-wise radiance field reconstruction, initializing each frame with the previous frame for temporal consistency. Figure 8 shows an example of synthesized test views for a motion clip. Notably, the pants are a non-rigid object that cannot be easily tracked, but GigaHands provides ample camera views for capturing non-rigid objects (more results in supp. document).
6 Conclusion
We present GigaHands, a massive annotated dataset of bimanual hand activities. The dataset contains 14k motion clips with 3D hand and object motions from 56 subjects interacting with 417 real-world objects, all captured from 51 camera views. It provides 183 million unique image frames and 84k textual descriptions, enabling a wide range of applications. We demonstrated how the scale and diversity of the dataset benefit text-driven motion synthesis and motion captioning, both within the dataset and on other data. Additionally, we have shown that GigaHands enables dynamic radiance field reconstruction, opening possibilities for downstream tasks.
Limitations and Future Directions. Despite its scale and diversity, GigaHands has limitations. The studio setting confines data collection to a limited space, making it challenging to accurately capture motions that require larger environments. While we can track rigid objects and object parts, fully automatic tracking of articulated and non-rigid objects remains challenging. Moreover, although we have showcased applications in motion synthesis and captioning, further research could explore how the dataset’s scale and diversity can enhance robotic manipulation and human-computer interaction tasks.
Acknowledgement
This research was supported by AFOSR grant FA9550-21-1-0214, NSF CAREER grant #2143576, and ONR DURIP grant N00014-23-1-2804. We would like to thank the OpenAI Research Access Program for API support and extend our gratitude to Ellie Pavlick, Tianran Zhang, Carmen Yu, Angela Xing, Chandradeep Pokhariya, Sudarshan Harithas, Hongyu Li, Chaerin Min, Xindi Qu, Xiaoquan Liu, Hao Sun, Melvin He and Brandon Woodard.